# Multipoint Schwarz–Pick Lemma for the quaternionic case

### v3

@@ 5.4. The case of nodes not in the real line
and the last term evaluates as $M_{p_{1}}(\tilde{p}_{\ell})^{-1}M_{s_{1}}(f(\hat{p}_{\ell}))$ where $\tilde{p}_{\ell}$ is a point in $\mathbb{S}_{p_{\ell}}$ depending on $p_{1}$ and $\hat{p}_{\ell}\in\mathbb{S}_{p_{\ell}}$ may depend on $p_{1},s_{1}$ and on the conjugate function $f^{c}$. They trivially coincide with $p_{\ell}$ when that point is real. So, the proposed algorithm does not work anymore, as $\hat{p}_{\ell}$ generally depends on the function $f^{c}$ which is not known a priori and in any case $f(\hat{p}_{\ell})$ is not known if $\hat{p}_{\ell}$ is not one of the nodes.
However, there is a specific situation in which we can recover a partial result in the spirit of Theorem 5.9. As Theorem 5.5 holds in the complex setting ([6, Thm. 4.1]) without the restriction to real nodes of interpolation, then we have the following partial answer: if $p_{1},\ldots,p_{n}\in\mathbb{B}$ belong to a common slice $\mathbb{C}_{I}$ and also $s_{1},\ldots,s_{n}\in\mathbb{B}\cap\mathbb{C}_{I}$, then the interpolation problem $f(p_{m})=s_{m}$, $m\in\{1,\ldots,n\}$ admits infinite solutions that are $\mathbb{C}_{I}$-preserving if $|Q_{n-1}^{n}|<1$, while it admits at least a solution given by a regular Blaschke product of degree at most $n-1$, when $|Q_{n-1}^{n}|=1$, and this is the unique one preserving $\mathbb{C}_{I}$.
@@
Since $q_{1},\ldots,q_{\kappa}$ belong to the same slice $\mathbb{C}_{I}$, then the regular Blaschke product $B=\mathcal{M}_{q_{1}}\ast\ldots\ast\mathcal{M}_{q_{\kappa}}v$ is such that $B\mid_{\mathbb{C}_{I}}=f_{0}$: hence the unique extension of $f_{0}$ to a slice regular self-map of the unit ball $\mathbb{B}$ is exactly $B$ and we get that a solution of the interpolation problem is given by a regular Blaschke product of degree less or equal than $n-1$, which preserves the slice $\mathbb{C}_{I}$. Any solution preserving that slice must coincide with $B$ on the slice and hence everywhere.

### v4

@@ 5.4. The case of nodes not in the real line
[…]
However, there is a specific situation in which we can recover a partial result in the spirit of Theorem 5.9. As Theorem 5.5 holds in the complex setting ([6, Thm. 4.1]) without the restriction to real nodes of interpolation, then we have the following partial answer: if $p_{1},\ldots,p_{n}\in\mathbb{B}$ belong to a common slice $\mathbb{C}_{I}$ and also $s_{1},\ldots,s_{n}\in\mathbb{B}\cap\mathbb{C}_{I}$, then the interpolation problem $f(p_{m})=s_{m}$, $m\in\{1,\ldots,n\}$ admits infinite solutions that are $\mathbb{C}_{I}$-preserving if $|Q_{n-1}^{n}|<1$, while it admits at least a solution given by a regular Blaschke product of degree at most n1n-1, when $|Q_{n-1}^{n}|=1$, and this is the unique one preserving $\mathbb{C}_{I}$.
@@
Since $q_{1},\ldots,q_{\kappa}$ belong to the same slice $\mathbb{C}_{I}$, then the regular Blaschke product $B=\mathcal{M}_{q_{1}}\ast\ldots\ast\mathcal{M}_{q_{\kappa}}v$ is such that $B\mid_{\mathbb{C}_{I}}=f_{0}$: hence the unique extension of $f_{0}$ to a slice regular self-map of the unit ball $\mathbb{B}$ is exactly $B$ and we get that a solution of the interpolation problem is given by a regular Blaschke product of degree less or equal than n1n-1, which preserves the slice $\mathbb{C}_{I}$. Any solution preserving that slice must coincide with $B$ on the slice and hence everywhere.